OpenStudy (anonymous):

Analyze the sequence below and complete the instructions that follow. 2, 4, 16, ... Use inductive reasoning to predict the next term in the sequence. A. 4 B. 32 C. 64 D. 256 Please select the best answer from the choices provided

OpenStudy (pratyush5):

If you observe then there is a pattern. The next number is square of present number

OpenStudy (pratyush5):

2^2 = 4 and 4^2 = 16

OpenStudy (pratyush5):

So you can find the next term

OpenStudy (anonymous):

ahh....idk

OpenStudy (pratyush5):

Next term is 16^2 = 16 X 16 . I assume you know basic multiplication ?